### v2

# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

Abstract
Weight sharing is a fundamental concept in neural architecture search (NAS), enabling gradient-based methods to explore cell-based architectural spaces significantly faster than traditional black-box approaches. In parallel, weight-entanglement has emerged as a technique for more intricate parameter sharing amongst macro-architectural spaces. Since weight-entanglement is not directly compatible with gradient-based NAS methods, these two paradigms have largely developed independently in parallel sub-communities. This paper aims to bridge the gap between these sub-communities by proposing a novel scheme to adapt gradient-based methods for weight-entangled spaces. This enables us to conduct an in-depth comparative assessment and analysis of the performance of gradient-based NAS in weight-entangled search spaces. Our findings reveal that this integration of weight-entanglement and gradient-based NAS brings forth the various benefits of gradient-based methods, while preserving the memory efficiency of weight-entangled spaces. The code for our work is openly accessible here.
1 Introduction
The concept of weight-sharing in Neural Architecture Search (NAS) was motivated by the need to improve efficiency over that of conventional black-box NAS algorithms, which demand significant computational resources to evaluate individual architectures. Here, weight-sharing (WS) refers to the paradigm by which we represent the search space with a single large supernet, also known as the one-shot model, that subsumes all the candidate architectures in that space. Every edge of this supernet holds all the possible operations that can be assigned to that edge. Importantly, architectures that share a particular operation also share its corresponding operation weights, allowing for efficient simultaneous partial training of an exponential number of subnetworks with each gradient update.
Gradient-based NAS algorithms (or optimizers), such as DARTS (liu-iclr19a), GDAS (dong-cvpr19a) and DrNAS (chen-iclr21a), assign an architectural parameter to every choice of operation on a given edge of the supernet. The output feature maps of these edges are thus an aggregation of the outputs of the individual operations on that edge, weighted by their architectural parameters. These architectural parameters are learned using gradient updates by differentiating through the validation loss. Supernet weights and architecture parameters are therefore trained simultaneously in a bi-level fashion. Once this training phase is complete, the final architecture can be obtained quickly, e.g., by selecting operations with the highest architectural weights on each edge as depicted in Figure 1(b). However, more sophisticated methods have also been explored (wang-iclr21a) for this selection.
While gradient-based NAS methods have primarily been studied for cell-based search spaces, a different class of search spaces focuses on macro-level structural decisions, such as the number of channels in a layer of the supernet, or the number of layers which are stacked to form the supernet. In these spaces, all architectures are subnetworks of the architecture with the largest architectural choices, identical to the supernet in this case. These search spaces share weights more intricately via weight-entanglement (WE) between similar operations on the same edge. An example of this for convolutional layers is that the 9 weights of a 3 $\times$ 3 convolution are a subset of the 25 weights of a 5 $\times$ 5 convolution. This paradigm reduces the memory footprint of the supernet to the size of the largest architecture in the space, unlike WS search spaces, where it increases linearly with the number of operation choices.
In order to efficiently search over such weight-entanglement spaces, two-stage methods first pre-train the supernet and then perform black-box search on it to obtain the final architecture. OFA (cai-iclr20a), SPOS (guo-eccv20a), AutoFormer (chen-iccv21b) and HAT (wang-acl20) are prominent examples of methods that fall into this category. Note that these methods do not employ additional architectural parameters for supernet training or search. They typically train the supernet by randomly sampling subnetworks and training them as depicted in Figure 1(a). The post-hoc black-box search relies on using the performance of subnetworks sampled from the trained supernet as a proxy for true performance on the unseen test set. To contrast with this two-stage approach, we refer to traditional gradient-based NAS approaches as single-stage methods.
Figure 1: (a) Two-Stage NAS with WE (Algorithm 3): dotted paths show operation choices not sampled at the given step (b) Single-Stage NAS with WS (Algorithm 4): every operation choice is evaluated independently and contributes to the output feature map with corresponding architecture parameters (c) Single-Stage NAS with WE (Algorithm 1): operation choices superimposed with corresponding architecture parameters. The architecture parameters for the three operation choices are represented by ${[\alpha_{i}]}_{i=1}^{3}$ and ${[\beta_{i}]}_{i=1}^{3}$. The operation weights, or choices, are symbolized by cubes (for convolutional layers) or rectangles (for feedforward layers) in various colors. In scenarios (a) and (c), due to weight entanglement, the smaller weights are effectively structured subsets of the larger weights. Conversely, in (b), through weight-sharing, operation weights are maintained independently from each other. In both (b) and (c), to determine the optimal architecture, the operations associated with the highest architecture parameter value are selected. This selection process applies to the choice of kernel size and the output dimension of the feedforward network.
To date, weight-entangled spaces have only been explored with two-stage methods, and cell-based spaces only with single-stage approaches. In this paper, we bridge the gap between these parallel sub-communities. We do so by addressing the challenges associated with integrating off-the-shelf single-stage NAS methods with weight-entangled search spaces. After a discussion of related work (Section 2), we make the following main contributions:
  1. 1.
    We propose a generalized scheme to apply single-stage methods to weight-entangled spaces while maintaining search efficiency and efficacy at larger scales (Section 3, with visualizations in Figure 1(c) and Figure 2 and Figure 3). We refer to this method as TangleNAS.
  2. 2.
    We propose a unified evaluation framework for the comparative evaluation of single and two-stage methods (Section 4.1) and study the effect of weight-entanglement in conventional cell-based search spaces (i.e., NAS-Bench-201 and the DARTS search space) (Section 4.2).
  3. 3.
    We evaluate our proposed generalized scheme for single-stage methods across a diverse set of weight-entangled macro search spaces and tasks, from image classification (Section 4.3.1 and Section 4.3.2) to language modeling (Section 4.3.3).
  4. 4.
    We conduct a comprehensive evaluation of the properties of single and two-stage approaches (Section 5), demonstrating that our generalized gradient-based NAS method achieves the best of single and two-stage methods: the enhanced performance, improved supernet fine-tuning properties, superior any-time performance of single-stage methods, and low memory consumption of two-stage methods. To facilitate reproducibility, our code is openly accessible here.
2 Related Work
[…]
DARTS, however, had prominent failure modes, such as its discretization gap and convergence towards parameter-free operations (white-arxiv23a), as outlined in Robust-DARTS (zela-iclr20a). Numerous gradient-based one-shot optimization techniques were developed since then (cai-iclr19a; nayman-neurips19a; wang-ijcai21; dong-cvpr19a; hu-cvpr20a; li-iclr21b; chen-iclr21a; zhang-icml21a). Amongst these, we highlight DrNAS (chen-iclr21a), which we will use in our experiments as a representative of gradient-based NAS methods. DrNAS treats one-shot search as a distribution learning problem, where the parameters of a Dirichlet distribution over architectural parameters are learned to identify promising regions of the search space. Despite the remarkable performance of single-stage methods, they are not directly applicable to some real-world architectural domains, such as transformers, due to the macro-level structure of these search spaces. DASH (shen-neurips22) employs a DARTS-like method to optimize CNN topologies (i.e., kernel size, dilation) for a diverse set of tasks, reducing computational complexity by appropriately padding and summing kernels with different sizes and dilations. FBNet-v2 (wan-cvpr20) and MergeNAS (wang-ijcai21) make an attempt along these lines for CNN topologies, but their methodology is not easily extendable to search spaces like transformers with multiple interacting modalities, such as embedding dimension, number of heads, expansion ratio, and depth.
Weight-entanglement, on the other hand, provides a more effective way of weight-sharing, exclusive to macro-level architectural spaces. In weight-entanglement, operations with weights of smaller dimensionality are structured subsets of the largest dimension. Hence, the total number of parameters in the supernet is the same as the number of parameters in the largest architecture in the space. The concept of weight-entanglement was developed in slimmable networks (yu-iclr18; yu-cvpr19), OFA (cai-iclr20a) and BigNAS (yu-eccv20) in the context of convolutional networks (see also AtomNAS (mei-iclr19)) and later spelled out in AutoFormer (chen-iccv21b) and applied to the transformer architecture.
Single-path-one-shot (SPOS) methods (guo-eccv20a) have shown a lot of promise in searching weight-entangled spaces. SPOS trains a supernet by uniformly sampling paths (one at a time to limit memory consumption) and then training the weights along that path. The supernet training is followed by a black-box search that uses the performance of the models sampled from the trained supernet as a proxy. OFA used a similar idea to optimize different dimensions of CNN architectures, such as its depth, width, kernel size, and resolution. Additionally, it enforced training of larger to smaller subnetworks sequentially to prevent interference between subnetworks. Subsequently, AutoFormer adopted the SPOS method to optimize a weight-entangled space of transformer architectures.
In this work, we demonstrate the application of single-stage methods to macro-level search spaces with weight-entanglement. This approach leverages the time efficiency and effectiveness of modern differentiable NAS optimizers, while maintaining the memory efficiency inherent to the weight-entangled space. Although DrNAS is our primary method for exploring weight-entangled spaces, our methodology is broadly applicable to other gradient-based NAS methods.
3 Methodology: Single-Stage NAS with Weight-Superposition
Figure 2: Weight superposition with architecture parameters ${\alpha_{i}}_{i=1}^{3}$ for kernel size search. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
Computational Efficiency
Our primary goal in this work is to effectively apply single-stage NAS to search spaces with macro-level architectural choices. To reduce the memory consumption and preserve computational efficiency, we propose two major modifications to single-stage methods. Firstly, the weights of the operations on every edge are shared with the corresponding weights of the largest operator on that edge. This reduces the size of the supernet to the size of the largest individual architecture in the search space.
Figure 3: Combi-superposition with parameters $\alpha_{i}\beta_{j}$. Supernet weight matrix (LHS) is adapted to gradient-based methods (RHS).
Secondly, to compute operation mixture, we take each operation choice, zero-pad to match the size of the largest choice, and then sum these, with each one multiplied by its respective architectural parameter. The resulting operation thus represents a mixture of different choices on an edge. This approach, visualized in Figure 1(c), contrasts with single-stage NAS methods, which weigh the outputs of operations on a specific edge using architectural parameters (refer to Figure 1(b)). Figure 2 provides an overview of the idea for a single architectural choice, such as the kernel size. This is equivalent to taking the largest operation and re-scaling the weights of each sub-operation by corresponding architectural parameters followed by summation of the weights (see the right-most weight matrix in Figure 2).
Algorithm 1 TangleNAS
1:Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$    $\mathcal{W}_{\max}\leftarrow\cup_{i-1}^{N}w_{i}$    $\mathcal{A}\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},...\alpha_{N}]$    $\gamma=$ learning rate of $\mathcal{A}$, $\eta=$ learning rate of $\mathcal{W}_{\max}$   $f$ is a function or distribution s.t. $\sum_{i=1}^{N}f(\alpha_{i})=1$
[…]
4:/* example of forward propagation on the cell */
5:for $j\leftarrow 1$ to $M$ do
6:   /* PAD weight to output dimension of $\mathcal{W}_{\max}$ before summation */
7:   /* Generalized Weighing Scheme */
8:   $\overline{o_{j}(x,\mathcal{W}_{\max})}=o_{j,i}(x,\sum_{i=1}^{N}f(\alpha_{i})\,\mathcal{W}_{\max}[:i])$
9:end for
10: /* weights and architecture update */
11: $\mathcal{A}=\mathcal{A}-\gamma\nabla_{\mathcal{A}}\mathcal{L}_{val}({\mathcal{W}_{\max}}^{*},\mathcal{A})$
12: $\mathcal{\mathcal{}}{W}_{\max}=\mathcal{W}_{\max}-\eta\nabla_{\mathcal{W}_{\max}}\mathcal{L}_{train}(\mathcal{W}_{\max},\mathcal{A})$
13: /* Architecture Selection */
14: $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
Weight Superposition
Weight Superposition is defined as a weighted summation of subsets of the largest weight matrix, to obtain weights with structural properties comparable to the largest operation. Consequently, a single forward pass on the superimposed weights suffices to capture the effect of all operational choices, thus making our approach computationally efficient. See Figure 2 for details on weight superposition of the kernel size.
Combi-Superposition
An operation may depend on two or more architectural decisions. Consider, for example, embedding dimension and intermediate MLP dimension (where the latter depends on the former by a searchable multiplicative factor i.e., the MLP ratio). To accommodate this, we introduce combi-superposition outlined in Figure 3 and in Algorithm 2. Combi-superposition superimposes the weights across multiple different architectural dimensions, allowing for search across different interacting architectural modalities.
Algorithm
These concepts allow us to apply any arbitrary gradient-based NAS method, such as DARTS, GDAS, or DrNAS, to macro-level search spaces that leverage weight-entanglement without incurring additional memory and computational costs during forward propagation. We name this single-stage architecture search method TangleNAS. See Algorithm 1 for an overview of the bi-level optimization framework with weight-superposition. The operation $f$ in Algorithm 1 determines the differentiable optimizer used in the method. For example, $f$ is a softmax function for DARTS and a function that samples from the Dirichlet distribution for DrNAS. We use DrNAS as the primary gradient-based NAS method in all our experiments, and refer to DrNAS used in conjunction with weight-entangled spaces as TangleNAS in the remainder of the paper.
4 Experiments
[…]
4.1 Toy search spaces
We begin the evaluation of TangleNAS on two compact toy search spaces that we designed as a contribution to the community to allow faster iterations of algorithm development:
  • Toy cell space: a small version of the DARTS space; architectures are evaluated on the Fashion-MNIST dataset (xiao2017fashion).
  • Toy conv-macro space: a small space inspired by MobileNet, including kernel sizes and the number of channels in each convolution layer as architectural decisions. The architectures are evaluated on CIFAR-10.
We describe these spaces in Appendix D, including links to code for these open source toy benchmarks. The results of these experiments are summarized in Tables 2 and 2. In both of these search spaces, TangleNAS outperforms its two-stage counterparts over 4 seeds. Additionally, DrNAS without weight-entanglement performs extremely poorly on the macro level search space (equivalent to a random classifier).
Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
Single-Stage DrNAS WS $91.190\pm 0.049$ 6.3
TangleNAS WE $\mathbf{91.300\pm 0.023}$ 6.2
Two-Stage SPOS+RS WE $90.680\pm 0.253$ 15.6
WE $90.317\pm 0.223$ 13.2
Optimum - - 91.630 -
Table 1: Evaluation on the toy cell-based search space on the Fashion-MNIST dataset.
Search Type Optimizer Supernet type Test acc (%) Search Time (hrs)
Single-Stage DrNAS WS $10\pm 0.000$ 12.4
TangleNAS WE $\mathbf{82.495\pm 0.461}$ 8.6
Two-Stage SPOS+RS WE $81.253\pm 0.672$ 21.7
WE $81.890\pm 0.800$ 26.4
Optimum - - 84.410 -
Table 2: Evaluation on the toy conv-macro search space on the CIFAR-10 dataset.
4.2 Cell-based search spaces
We now begin our comparative analysis of single and two-stage approaches by applying them to cell-based spaces, which are central in the single-stage NAS literature. We evaluate TangleNAS against DrNAS and SPOS on these spaces. Here, we use the widely studied NAS-Bench-201 (NB201) (dong2020bench) and DARTS (liu-iclr19a) search spaces. We refer the reader to Appendix D for details about these spaces and Appendix F.4 for the experimental setup. We evaluate each method with 4 different random seeds.
Search Type Optimizer Supernet CIFAR-10 (%) ImageNet (%) Search Time (hrs)
Single-Stage DrNAS WS $2.625\pm 0.075$ 26.290 9.1
TangleNAS WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.691}$ 7.4
Two-Stage SPOS+RS WE $2.965\pm 0.072$ 27.114 18.7
SPOS+ES WE $3.200\pm 0.065$ 27.320 14.8
Table 3: Comparison of test errors of single and two-stage methods on the DARTS search space.
Our contribution on these spaces is two-fold. Firstly, we study the effects of weight-entanglement on cell-based spaces in conjunction with single-stage methods. To this end, we entangle the weights of similar operations with different kernel sizes on both search spaces. For NB201, the weights of the $\times1$ and $\times3$ convolutions are entangled, and in the DARTS search space, the weights of dilated and separable convolutions with kernel sizes $\times3$ and $\times5$ are (separately) entangled. Secondly, we study SPOS on the NB201 and DARTS search spaces. To the best of our knowledge, we are the first to study a two-stage method like SPOS in such cell search spaces.
Tables 3 and 4 show the results. For both search spaces, TangleNAS yields the best results, outperforming the single-stage baseline DrNAS with WS, as well as both SPOS variants. TangleNAS also significantly lowers the memory requirements and runtime compared to its weight-sharing counterpart. We note that overall, the SPOS methods are ineffective in these cell search spaces.
Search Type Optimizer Supernet CIFAR-10 CIFAR-100 ImageNet16-120 Search Time (hrs)
Single-Stage DrNAS WS $\mathbf{94.360\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.370\pm 0.000}$ 20.9
TangleNAS WE $\mathbf{94.360\pm 0.000}$ $\mathbf{73.510\pm 0.000}$ $\mathbf{46.370\pm 0.000}$ 20.4
Two-Stage SPOS+RS WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$ 29.5
SPOS+ES WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$ 26.7
Table 4: Comparison of test-accuracies of single and two-stage methods on NB201 search space.
4.3 Macro Search Spaces
Given the promising results of TangleNAS on the toy and cell-based spaces, we now extend our evaluation to the home base of two-stage methods. We study TangleNAS on a vision transformer space (AutoFormer-T and -S) and a convolutional space (MobileNetV3), both of which were previously proposed and examined using two-stage methods by chen-iccv21b and cai-iclr20a, respectively. Additionally, we explore a language model transformer search space centered around GPT-2 (radford-openaiblog19a).
4.3.1 AutoFormer
Search Type Optimizer CIFAR-10 CIFAR-100
Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS Inherit (%) Fine-tune (%) Retrain (%) Params FLOPS
Single-Stage TangleNAS 93.570 $\mathbf{97.702\pm 0.017}$ $\mathbf{97.872\pm 0.054}$ 8.685M 1.946G 76.590 $\mathbf{82.615\pm 0.064}$ $\mathbf{82.668\pm 0.161}$ 8.649M 1.939G
Two-Stage SPOS+RS 94.290 $97.605\pm 0.038$ $97.767\pm 0.024$ 8.512M 1.910G 78.210 $82.407\pm 0.026$ $82.210\pm 0.142$ 8.476M 1.905G
SPOS+ES 94.100 $97.632\pm 0.047$ $97.643\pm 0.023$ 7.230M 1.659G 77.970 $82.517\pm 0.140$ $82.518\pm 0.114$ 8.245M 1.859G
Table 5: Test Accuracies on the AutoFormer-T space for CIFAR-10 and CIFAR-100 (across 4 random seeds)
We evaluate TangleNAS on the AutoFormer-T and AutoFormer-S spaces introduced by chen-iccv21b, based on vision transformers. The search space consists of the choices of embedding dimensions and number of layers, and for each layer, its MLP expansion ratio and the number of heads it uses to compute attention. More details can be found in Table 20 in the appendix. The embedding dimension remains constant across the network, while the number of heads and the MLP expansion ratio change for each layer. This results in a search space of about $10^{13}$ architectures. We train our supernet using the same training hyperparameters and pipeline as used in AutoFormer, and use its evolutionary search as our baseline.
SuperNet-Type NAS Method ImageNet (%) Datasets Params FLOPS
CIFAR-10 (%) CIFAR-100 (%) Flowers (%) Pets(%) Cars(%)
AutoFormer-T SPOS+ES 75.474 98.019 86.369 98.066 91.558 91.935 5.893M 1.396G
AutoFormer-T TangleNAS 78.842 98.249 88.290 98.066 92.347 92.396 8.981M 2.000G
AutoFormer-S SPOS+ES 81.700 99.100 90.459 97.900 94.853 92.545 22.900M 5.100G
AutoFormer-S TangleNAS 81.964 99.120 90.459 98.326 95.070 92.371 28.806M 6.019G
Table 6: Evaluation on the AutoFormer-T space on downstream tasks. ImageNet-1k validation accuracies are obtained through inheritance, whereas the test accuracies for the other datasets are achieved through fine-tuning the ImageNet-pretrained model.
In Tables 5 and 6, we evaluate TangleNAS against AutoFormer on the AutoFormer-T and AutoFormer-S spaces. Interestingly, we observe that although AutoFormer sometimes outperforms TangleNAS upon inheritance from the supernet, the TangleNAS architectures are always better upon fine-tuning and much better upon retraining. For ImageNet-1k we obtain an improvement of $3.368\%$ and $0.264\%$ on AutoFormer-T and AutoFormer-S spaces, respectively (see Table 6).
4.3.2 MobileNetV3
Next, we study a convolutional search space based on the MobileNetV3 architecture. The search space is defined in Table 19 in the appendix and contains about $2\times 10^{19}$ architectures. This follows from the search space designed by OFA (cai-iclr20a), which searches for kernel size, number of blocks, and channel-expansion factor.
Search Type Optimizer Top-1 acc (%) Params FLOPS
Single-Stage TangleNAS 77.424 7.580M 528.80M
Two-Stage OFA+RS 77.046 6.870M 369.160M
OFA+ES 77.050 7.210M 420.500M
Largest Arch - 77.336 7.660M 566.170M
Table 7: Evaluation on MobileNetV3.
During the supernet training, we activate all choices in our supernet at all times. Table 7 shows that on this OFA search space, TangleNAS outperforms OFA (based on both unconstrained evolutionary and random search on the pre-trained OFA supernet). Notably, TangleNAS even yields an architecture with higher accuracy than the largest architecture in the space (while OFA yields worse architectures).
4.3.3 Language Modeling (LM) Space
Finally, given the growing interest in efficient large language models and recent developments in scaling laws (kaplan2020scaling; hoffmann2022training), we study our efficient and scalable TangleNAS method on a language-model space for two different scales. We create our language model space around a smaller version of nanoGPT model [^1] https://github.com/karpathy/nanoGPT and the model at its original size. In this transformer search space, we search (with 4 random seeds) for the embedding dimension, number of heads, number of layers, and MLP ratios (as defined in Table 18 of the appendix). We weight-entangle all of these by combi-superposition in four dimensions.
Architecture Search-Type Loss Perplexity $\mathbf{\downarrow}$ Params Inference time (s)
GPT-2 Manual 3.077 21.690 123.590M 113.30
TangleNAS Automated 2.904 18.243 116.519M 102.50
(a) Comparison of fine-tuning architecture discovered by TangleNAS and GPT-2 on Shakespeare dataset.
Search Type Optimizer Test loss Perplexity $\mathbf{\downarrow}$ Params Inference Time (s)
Single-Stage TangleNAS $\mathbf{1.412\pm 0.011}$ 4.104 87.010M 93.88
Two-Stage SPOS+RS $1.433\pm 0.005$ 4.191 77.570M 85.17
SPOS+ES $1.444\pm 0.013$ 4.238 78.750M 87.10
(b) Comparison of single and two-stage methods on language model search space. We report the test loss and perplexity on the TinyStories dataset.
Table 8: Evaluation on Language Model Space
For each of these transformer search dimensions, we consider 3 different choices.We train our language models on the TinyStories (eldan-arxiv2023) dataset (for the smaller version of nanoGPT) and on OpenWebText (Gokaslan2019OpenWeb) (for nanoGPT at its original size). Furthermore, we fine-tune the model trained on OpenWebText on the Shakespeare dataset. On the smaller scale, we beat the SPOS+ES and SPOS+RS baselines as presented in Table 8(b). This improvement is statistically very significant with a two-tailed p-value of $0.0064$ for ours v/s SPOS+ES and p-value of $0.0127$ for ours v/s SPOS+RS. As shown in Table 8(a), we obtain a smaller model on OpenWebText while achieving better perplexity after fine-tuning on Shakespeare [^2] https://huggingface.co/datasets/karpathy/tiny_shakespeare. This model is also very efficient during inference time in comparison to GPT-2.
5 Results and Discussion
For a more thorough evaluation, we now compare different properties of single and two-stage methods, focusing on their any-time performance, the impact of the train-validation split ratios, and the Centered Kernel Alignment (kornblith2019similarity) (see Appendix H) of the supernet feature maps. Additionally, we study the effect of pretraining, fine-tuning and retraining on the AutoFormer space for CIFAR-10 and CIFAR-100, as well as the downstream performance of the best model pre-trained on ImageNet across various classification datasets. We conclude by discussing the insights derived from TangleNAS in designing architectures on real world tasks.
Space and time complexity
In practice, we observe that vanilla gradient-based NAS methods are memory and compute expensive in comparison to both two-stage methods and our TangleNAS approach with weight-superposition. While the time and space complexity of single-stage methods is $\mathcal{O}(n)$, where $n$ is the number of operation choices, TangleNAS, similar to two-stage methods, maintains $\mathcal{O}(1)$. On the NB201 and DARTS search spaces we observe a 25.28% and 35.54% reduction in memory requirements for TangleNAS over DrNAS with WS. This issue only exacerbates for weight-entangled spaces like AutoFormer, MobileNetV3 and GPT, making the application of vanilla gradient-based methods practically infeasible.
Anytime performance
NAS practitioners often emphasize rapid discovery of competitive architectures. This is especially important given the rising costs of training large and complex neural networks, like Transformers. Thus, strong anytime performance is crucial for the practical deployment of NAS in resource-intensive environments. Therefore, we examine the anytime performance of TangleNAS. Figure 4 demonstrates that TangleNAS (DrNAS+WE) is faster than its baseline method (DrNAS+WS). Similarly, Figure 5 shows that TangleNAS has better anytime performance than AutoFormer.
(a) CIFAR-100
(b) ImageNet16-120
Figure 4: Test accuracy evolution over epochs for NB201.
(a) CIFAR-10
(b) CIFAR-100
Figure 5: Any Time performance curves of AutoFormer vs. Ours.
Effect of fraction of training data
One-shot NAS commonly employs a 50%-50% train-valid split for cell-based spaces and 80%-20% for weight-entangled spaces. To eliminate possible biases, we tested our method on various data splits within each search space. The findings, detailed in Section B, show consistent results across these splits. Specifically, single-stage methods prove to be robust and performant across different training fractions and search spaces, compared to two-stage methods.
Figure 6: MLP ratio trajectory for LM. Number of layers range from 1-12 and MLP ratio choice can be 2, 3 or 4.
5.1 Insights from NAS
Architecture design insights
In transformer spaces, reducing the MLP-ratio in the initial layers has a relatively low impact on performance (Figure 6) and can often work competitively or outperform handcrafted architectures. This observation is consistent across ViT and Language Model spaces. Conversely, number of heads and embedding dimension have a significant impact.
Pruning a few of the final layers also has a relatively low impact on performance. In the MobileNetV3 space, we find a strong preference for a larger number of channels and larger network depth. In contrast, we discover that scaling laws for transformers may not necessarily apply in convolutional spaces, especially for kernel sizes - the earlier layers favor $\times5$ kernel sizes while later ones prefer $\times7$ (3 being the smallest and 7 the largest).
Effect of pretraining, fine-tuning and retraining
We examined the effects of inheriting, fine-tuning, and retraining in the AutoFormer space on the CIFAR-10 and CIFAR-100 datasets. We observe that retraining generally surpasses both fine-tuning and inheriting. This raises questions about the correlation between inherited and retrained accuracies of architectures in two-stage methods and the potential training interference identified by xu2022analyzing. A strong correlation is crucial for two-stage methods, which use the performance of architectures with inherited weights as a proxy for true performance in the black-box search. Indeed, while the SPOS+RS and SPOS+ES methods perform well with inherited weights, TangleNAS exceeds their performance after fine-tuning and retraining.
ImageNet-1k pre-trained architecture on downstream tasks
Lastly, we study the impact on fine-tuning the best model obtained from the search on downstream datasets. We follow the fine-tuning pipeline proposed in AutoFormer and fine-tune on different fine- and coarse-grained datasets. We observe from Table 6 that the architecture discovered by TangleNAS on ImageNet is much more performant in fine-tuning to various datasets (CIFAR-10, CIFAR-100, Flowers, Pets and Cars) than the architecture discovered by SPOS.
6 Conclusion and Broader Impact
In this paper, we compare single-stage and two-stage NAS methods, traditionally used for different search spaces, and introduce single-stage NAS to weight-entangled spaces, usually the domain of two-stage methods. We empirically evaluate our single-stage method, TangleNAS, on a diverse set of weight-entangled search spaces and tasks, showcasing its ability to outperform conventional two-stage NAS methods while enhancing search efficiency. Our positive results on macro-level search spaces suggest this approach could advance the development of modern architectures like Transformers within the NAS community. A recent work (klein2023structural) starts training of the supernet from the largest pretrained model, subsequently fine-tuning it. Our method, which now renders single-stage methods applicable to broader families of search spaces (e.g., transformers), can similarly benefit from initialization with pretrained models, achieving additional computational savings. We leave this for future work.
This study addresses the high energy consumption associated with neural architecture search, particularly in black-box techniques that demand extensive computational resources to train many architectures from scratch. Our proposed method falls in the family of one-shot NAS methods, and hence significantly reduces energy usage and identifies efficient architectures with far less energy consumption than manual tuning.
Acknowledgments
This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Robert Bosch GmbH is acknowledged for financial support. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix A Submission Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    2. (b)
      Did you describe the limitations of your work? Yes, in Appendix I
    3. (c)
      Did you discuss any potential negative societal impacts of your work? NA
    4. (d)
      Did you read the ethics review guidelines and ensure that your paper conforms to them? https://2022.automl.cc/ethics-accessibility/ Yes
  2. 2.
    If you ran experiments…
    1. (a)
      Did you use the same evaluation protocol for all methods being compared (e.g., same benchmarks, data (sub)sets, available resources)? Yes
    2. (b)
      Did you specify all the necessary details of your evaluation (e.g., data splits, pre-processing, search spaces, hyperparameter tuning)? Yes
    3. (c)
      Did you repeat your experiments (e.g., across multiple random seeds or splits) to account for the impact of randomness in your methods or data? Yes, for experiments on smaller scales
    4. (d)
      Did you report the uncertainty of your results (e.g., the variance across random seeds or splits)? Yes, for experiments on smaller scales
    5. (e)
      Did you report the statistical significance of your results? Yes, we perform a two tailed p-value test on the tinystories dataset for language modeling
    6. (f)
      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes, We used NAS-Bench-201 which is tabular
    7. (g)
      Did you compare performance over time and describe how you selected the maximum duration? Yes, we study performance over compute budgets
    8. (h)
      Did you include the total amount of compute and the type of resources used (e.g., type of gpus, internal cluster, or cloud provider)? Yes
    9. (i)
      Did you run ablation studies to assess the impact of different components of your approach? No, We only add a single main component in our approach
  3. 3.
    With respect to the code used to obtain your results…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results, including all requirements (e.g., requirements.txt with explicit versions), random seeds, an instructive README with installation, and execution commands (either in the supplemental material or as a url)? Yes
    2. (b)
      Did you include a minimal example to replicate results on a small subset of the experiments or on toy data? Yes
    3. (c)
      Did you ensure sufficient code quality and documentation so that someone else can execute and understand your code? Yes
    4. (d)
      Did you include the raw results of running your experiments with the given code, data, and instructions? Yes, we include the smaller pickle files for toy benchmarks we create and results logs wherever possible. We will share larger raw files eg: checkpoints upon acceptance, as anonymization of these resources is difficult.
    5. (e)
      Did you include the code, additional data, and instructions needed to generate the figures and tables in your paper based on the raw results? Yes
  4. 4.
    If you used existing assets (e.g., code, data, models)…
    1. (a)
      Did you cite the creators of used assets? Yes
    2. (b)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating if the license requires it? We do not use personal data
    3. (c)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? We do not use personal data
  5. 5.
    If you created/released new assets (e.g., code, data, models)…
    1. (a)
      Did you mention the license of the new assets (e.g., as part of your code submission)? Yes
    2. (b)
      Did you include the new assets either in the supplemental material or as a url (to, e.g., GitHub or Hugging Face)? Yes
  6. 6.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? NA
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (irb) approvals, if applicable? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
  7. 7.
    If you included theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
Appendix B Training across data fractions
TangleNAS is more robust across dataset fractions for network weights and architecture optimization as seen from Tables 9, 10, 11, and 12.
Search Type Optimizer Train portion CIFAR-10 (%) CIFAR-100 (%)
Single-Stage TangleNAS 50% $97.715\pm 0.088$ $82.538\pm 0.118$
80% $\mathbf{97.872\pm 0.054}$ $\mathbf{82.668\pm 0.161}$
Two-Stage SPOS+RS 50% $97.680\pm 0.026$ $82.537\pm 0.280$
80% $97.767\pm 0.024$ $82.210\pm 0.142$
SPOS+ES 50% $97.77\pm 0.038$ $82.354\pm 0.120$
80% $97.642\pm 0.023$ $82.517\pm 0.114$
Table 9: Evaluation on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
Search Type Optimizer Train portion Supernet Accuracy (%)
CIFAR-10 CIFAR-100 ImageNet16-120
Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $71.153\pm 0.697$ $\mathbf{46.37\pm 0.00}$
TangleNAS 50% WE $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $\mathbf{73.51\pm 0.000}$ $\mathbf{46.37\pm 0.00}$
Two-Stage SPOS+RS 50% WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$
80% $87.778\pm 2.446$ $53.68\pm 4.174$ $30.545\pm 3.643$
SPOS+ES 50% WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$
80% $89.095\pm 0.825$ $56.363\pm 4.724$ $30.935\pm 3.546$
Table 10: Comparison of test accuracies of single and two-stage methods with WS and WE on NB201 search space.
Search Type Optimizer Train portion Supernet type Test acc (%)
Single-Stage DrNAS 50% WS $91.19\pm 0.049$
80% $91.125\pm 0.033$
TangleNAS 50% WE $\mathbf{91.3\pm 0.023}$
80% $91.065\pm 0.163$
Two-Stage SPOS+RS 50% WE $90.68\pm 0.253$
80% $90.687\pm 0.110$
SPOS+ES 50% WE $90.318\pm 0.223$
80% $90.595\pm 0.219$
Optimum - - - 91.63
Table 11: Evaluation on toy cell-based search space on Fashion-MNIST dataset.
Search Type Optimizer Train portion Supernet type Test acc (%)
Single-Stage DrNAS 50% WS 10
80% 10
TangleNAS 50% WE $\mathbf{83.020\pm 0.000}$
80%   $82.495\pm.0.461$
Two-Stage SPOS+RS 50% WE $81.253\pm 0.672$
80% $81.345\pm 0.383$
SPOS+ES 50% WE $81.890\pm 0.800$
80%    $82.322\pm 0.604$
Optimum - - - 84.41
Table 12: Evaluation on toy conv-macro search space on CIFAR-10 dataset.
Search Type Optimizer Train portion Supernet CIFAR-10 (%) ImageNet (%)
Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.69}$
80% $2.67\pm 0.076$ 25.742
DrNAS 50% WS $2.625\pm 0.075$ 26.29
80% $2.580\pm 0.028$ $\mathbf{25.67}$
Two-Stage SPOS+RS 50% WE $2.965\pm 0.072$ 27.114
80% $2.965\pm 0.072$ 27.114
SPOS+ES 50% WE $3.200\pm 0.066$ 27.424
80% $3.002\pm 0.037$ 26.76
Table 13: Comparison of test errors of single and two-stage methods with WS and WE on DARTS search space.
Appendix C Comparison with different gradient-based optimizers
We also compare DrNAS against different gradient-based NAS optimizers in Tables 15, 14, and 16 on the toy cell-based, toy macro and the AutoFormer search spaces. We observe that DrNAS outperforms GDAS and DARTS on all of these search spaces, showing its robust nature.
Optimizer Test-acc (%)
TangleNAS+DrNAS 83.02
SPOS+RS 81.25
SPOS+ES 81.89
TangleNAS+DARTS_v1 81.61
TangleNAS+DARTS_v2 81.49
TangleNAS+GDAS 10 (degenerate)
Table 14: Comparison of DrNAS with other gradient-based optimizers on the toy conv-macro search space on CIFAR-10.
Optimizer Test-acc (%)
TangleNAS+DrNAS 90.930
SPOS+RS 90.688
SPOS+ES 90.595
TangleNAS+DARTS_v1 89.905
TangleNAS+DARTS_v2 90.747
TangleNAS+GDAS 90.618
Table 15: Comparison of DrNAS with other gradient-based optimizers on the toy cell-based search space on FashionMNIST dataset.
Optimizer CIFAR-10 CIFAR-100
TangleNAS+DrNAS 97.872 $\pm$ 0.054 82.668 $\pm$ 0.161
SPOS+RS 97.767 $\pm$ 0.024 82.210 $\pm$ 0.142
SPOS+ES 82.518 $\pm$ 0.114 82.518 $\pm$ 0.114
TangleNAS+DARTS_v1 97.672 $\pm$ 0.040 82.107 $\pm$ 0.392
TangleNAS+GDAS 97.45 $\pm$ 0.096 82.120 $\pm$ 0.281
Table 16: Comparison of DrNAS with other gradient-based optimizers on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
Appendix D Search Space details
Toy cell space
This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable $\times3$ Convolution, (b) Separable $\times5$ Convolution, (c) Dilated $\times3$ Convolution, and (d) Dilated $\times5$ Convolution. The macro-architecture of the model comprises three cells of the types reduction, normal, and reduction again stacked together. Notably, we entangle the $\times3$ and $\times5$ kernel weights for each operation type, i.e., separable convolutions and dilated convolutions. We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark, which we release here.
Toy conv-macro space
This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also referred to as a macro search space) for four convolutional layers. Every convolutional layer has a choice of three kernel sizes and number of channels. See Table 17 for more details. We evaluate this search space on the CIFAR-10 dataset by creating a small benchmark, which we release here.
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Number of channels (layer 1) 8, 16, 32
Number of channels (layer 2) 16, 32, 64
Number of channels (layer 3) 32, 64, 128
Number of channels (layer 4) 64, 128, 256
Table 17: Toy Convolutional-Macro Search Space.
NAS-Bench-201 Space
The NAS-Bench-201 search space (dong2020bench) has a single cell type which is stacked 15 times, with residual blocks (he-cvpr16a) after the fifth and tenth cells. Each cell has 4 nodes, with each node connected to the previous ones by operations chosen from skip connection, $\times3$ convolution, $\times1$ convolution, $\times3$ average-pooling, and zeroize, which zeros out the input feature map. There is limited scope for weight-entanglement in this space, given that only two of the five candidate operations contain learnable parameters. In this case, we entangle the $\times3$ and $\times1$ convolution kernels on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
Architectural Parameter Choices
Embedding Dimension 384, 576, 768
Number of Heads 6, 8, 12
MLP Expansion Ratio 2, 3, 4
Number of Layers 5,6,7
Table 18: Choices for Language Model Space.
DARTS Search Space
The DARTS (liu-iclr19a) search space has two kinds of cells - normal and reduction. Reduction cells double the number of channels in its outputwhile halving the width and height of the feature maps. The supernet consists of 8 cells stacked sequentially, while the discretized network stacks 20 cells. Reduction cells are placed at 1/3 and 2/3 of the total depth of the network. Each cell takes two inputs - one each from the outputs of the two previous cells. There are 14 edges in each cell, with each edge in the supernet consisting of 8 candidate operations as follows: $\times3$ max pooling, $\times3$ average pooling, skip connect, $\times3$ separable convolutions, $\times5$ separable convolutions, $\times3$ dilated convolutions, $\times5$ dilated convolutions, and none (no operation).
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Channel expansion (all layers) 3,4,6
Number of blocks 2,3,4
Table 19: MobileNetV3 Search Space.
AutoFormer Space and Language Model Space
We present the details of the AutoFormer Space and the Language Model space is Tables 20 and 18, respectively.
Architectural Parameter Choices
Embedding dimension 192, 216, 240
Number of layers 12, 13, 14
MLP ratio (per layer) 3.5, 4
Number of heads (per layer) 3, 4
Table 20: Choices for AutoFormer-T Search Space.
Appendix E Methodological details
Combi-Superposition
Traditional cell-based search spaces primarily consider independent operations (e.g., convolution or skip). One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces, which are not general enough to be applied to macro-level architectural parameters. Consider, e.g., the task of searching for the embedding dimension and the expansion ratio for a transformer. Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters – one corresponding to the choice of embedding dimension and the other to the expansion ratio. To adapt single-stage methods to these combined operation choices in the search space, we propose the combi-superposition operation.
The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions. This allows us to optimize jointly in this space without the need for a separate forward pass for each combination. Every combination maps to a unique sub-matrix of the operator weight matrix, indexed by both the embedding dimension and the expansion ratio. To address shape mismatches of the different operation weights during forward passes, every sub-matrix is zero-padded to match the shape of the largest matrix. See Algorithm 2 for more details, and Figure 3 for an overview of the idea.
Algorithm 2 Combi-Superposition Operation
For ease of presentation, we show the algorithm for superimposing along two dimensions. However, in practice, we can superimpose an arbitrary number of dimensions (e.g., four dimensions in our experiments with NanoGPT).
$embed\_dim\leftarrow[e_{1},e_{2},e_{3},\ldots,e_{n}]$ {Choices for embedding dimension}
$expansion\_ratio\leftarrow[r_{1},r_{2},r_{3},\ldots,r_{m}]$ {Choices for expansion ratio}
$\alpha\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},\ldots,\alpha_{n}]$ {Architecture parameters for embedding dimension}
$\beta\leftarrow[\beta_{1},\beta_{2},\beta_{3},\ldots,\beta_{m}]$ {Architecture parameters for expansion ratio}
$X\leftarrow input\_feature$
$W,b\leftarrow fc\_layer\_weight,fc\_layer\_bias$
$W_{mix}\leftarrow\textbf{0}$
$b_{mix}\leftarrow\textbf{0}$
for $i\leftarrow 1$ to $n$ do
  for $j\leftarrow 1$ to $m$ do
   $W\_{ij}=W[:(embed\_dim[i]\times expansion\_ratio[j]),:embed\_dim[i]]$
   $b\_{ij}=b[:embed\_dim[i]\times expansion\_ratio[j]]$
   $W_{\text{mix}}\leftarrow W_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(W_{ij})$
   $b_{\text{mix}}\leftarrow b_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(b_{ij})$
  end for
end for
$Y\leftarrow X\cdot W_{mix}+b_{mix}$ {Compute the output of the FC layer with a mixture of weights and bias}
return $Y$
For completeness and to facilitate comparison with TangleNAS (Algorithm 1), we present Algorithm 3, which describes a generic two-stage method on a macro search space with weight entanglement (WE). Additionally, vanilla single-stage methods on cell-based weight sharing (WS) spaces follow Algorithm 4.
Algorithm 3 Weight Entanglement (Two-Stage)
1:Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$   $\mathcal{W}_{\max}\leftarrow\cup_{i-1}^{N}w_{i}$    $\eta=$ learning rate of $\mathcal{W}_{\max_{\mathcal{O}}}$
[…]
4:/* example of forward propagation on the cell */
5:for $j\leftarrow 1$ to $M$ do
6:   $i\sim\mathcal{U}(1,N)$
7:   /* Slice weight matrix corresponding to operation */
8:   $o_{j}(x,\mathcal{W}_{max})=o_{(j,i)}(x,\mathcal{W}_{max}[:i])$
9:end for
10: /* weights update */
11: $\mathcal{W}_{\max}[:i]=\mathcal{W}_{\max}[:i]-\eta\nabla_{\mathcal{W}_{\max}}[:i]\mathcal{L}_{train}(\mathcal{W}_{\max})$
[…]
Algorithm 4 Weight Sharing (Single-Stage)
1:Input: $M\leftarrow$ number of cells, $N\leftarrow$ number of operations    $\mathcal{O}\leftarrow[o_{1},o_{2},o_{3},...o_{N}]$    $\mathcal{W_{O}}\leftarrow[w_{1},w_{2},w_{3},....w_{N}]$    $\mathcal{A}\leftarrow[\alpha_{1},\alpha_{2},\alpha_{3},...\alpha_{N}]$    $\gamma=$ learning rate of $\mathcal{A}$   $\eta=$ learning rate of $\mathcal{W_{O}}$    $f$ is a function or distribution s.t. $\sum_{i=1}^{N}f(\alpha_{i})=1$
[…]
4:/* example of forward propagation on the cell */
5:for $j\leftarrow 1$ to $M$ do
6:   /* Compute mixture operation as weighted sum of output of operations*/
7:   $\overline{o_{j}(x,\mathcal{W_{O}})}=\sum_{i=1}^{N}f(\alpha_{i})\,o_{(j,i)}(x,w_{(j,i)})$
8:end for
9: /* weights and architecture update */
10: $\mathcal{A}=\mathcal{A}-\gamma\nabla_{\mathcal{A}}\mathcal{L}_{val}({\mathcal{W_{O}}}^{*},\mathcal{A})$
11: $\mathcal{W_{O}}=\mathcal{W_{O}}-\eta\nabla_{\mathcal{W_{O}}}\mathcal{L}_{train}(\mathcal{W_{O}},\mathcal{A})$
12: /* Search */
13: $selected\_arch\leftarrow$ $\arg\max(\mathcal{A})$
Compatibility issues between weight-entanglement and gradient-based methods
We address the incompatibility between gradient-based NAS methods and weight-entangled (WE) spaces as follows (where n refers to the number of operation choices):
  • Gradient-based NAS methods do not share or entangle weights among competing operations, which increases their GPU memory footprint. We tackle this issue by adopting weight-entanglement from two-stage methods, thereby reducing the parameter size of the supernet from $\mathcal{O}$}(n) to $\mathcal{O}$}(1).
  • Gradient-based NAS methods compute a weighted combination of output of the operations. Even after entangling the operation weights, this approach leads to increased GPU memory usage because all intermediate competing activations/features need to be preserved in memory. Additionally, this method does not scale well with an increase in the number of operation choices, as GPU memory consumption during forward propagation scales as $\mathcal{O}$}(n). To address this, we propose weight superposition, which computes the weighted combination directly within the space of entangled weights.
  • In vanilla gradient-based methods, a forward pass is computed independently for each operation choice, resulting in a time and memory complexity of $\mathcal{O}$}(n). In contrast, TangleNAS computes only a single forward pass using the superimposed weights, reducing the cost of the forward pass to $\mathcal{O}$}(1).
Details on Figure 1
In the overview Figure 1, we use rounded squares to represent the input and output feature maps of a convolution. The colored cubes represent the convolutional kernels, while the colored rectangles denotes non-convolutional architectural choices. These two differ primarily in how their weights are entangled.
Consider Figure 1 (a), which illustrates the forward pass of an input feature map through the candidate operations on an edge of the supernet in a two-stage weight-entanglement method (such as OFA). In two-stage methods, random paths through the supernet are sampled and trained in the first stage. In this figure, thick lines indicate the paths sampled in a given step. The weights of the operation choices, depicted in the figure by colored cubes and rectangles, overlap with one another, showing that all the operations use slices of a common, large weight matrix. We show three choices of convolutions, each of a different kernel size (say, $\times1$, $\times3$, and $\times5$), represented by cubes of varying sizes. The $\times1$ and $\times3$ convolutions use slices of the larger $\times5$ convolution as their weights, represented in the figure by nesting the smaller kernels inside the larger ones. Since only one operation is sampled at a time (in this case, the orange kernel), there is only one output feature map. This feature map, after global average pooling, is then passed through one of the three non-convolutional operation choices. The operation choices here may be the embedding dimension, for example, and different slices of the largest feedforward network are used for the choices of embedding dimensions. At the end of the first stage in Figure 1 (a), the supernet has been trained along different paths. In the second stage, paths are sampled from the trained supernet using black-box methods (such as random or evolutionary search) and evaluated on the dataset to obtain the optimal architecture, which we represent with optim arch.
Now, consider Figure 1 (b), which represents the forward pass in a single-stage weight-sharing method (such as DARTS). As you can see, there is no nesting of the weights of the three convolutions or feedforward networks in this case, indicating that each has its own distinct set of weights. Naturally, this will incur more GPU memory usage, as more weights need to be stored and their gradients computed. The outputs from each of these convolutional (or feedforward) operations are then weighted by $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ (or $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$), which represent the architectural parameters of the operations. These weighted feature maps are then summed up to produce the output feature maps.
The optimization loop of single-stage weight-sharing methods, shown in Figure 1(b), aims to find the optimal values for the architectural parameters $\alpha$ and $\beta$. We obtain the discretized model by selecting only the operations in the supernet with the highest values of these parameters at the end of this loop. Specifically, the convolutional kernel and non-convolutional operator (in this case, the feedforward network) with the maximum values of $\alpha$ and $\beta$ are represented as argmax($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) and argmax($\beta_{1}$, $\beta_{2}$, $\beta_{3}$), respectively. The resulting discretized model, or optimal architecture, is denoted as optimal arch. Note that these methods do not require black-box search, as the optimal architecture can be obtained directly from the learned architectural parameters.
In Figure 1 (c), we depict our hybrid framework, which utilizes both entangled (nested) weights and architectural parameters. This approach reduces GPU memory usage due to weight entanglement and allows for the optimal architecture to be obtained directly from the learned parameters, eliminating the need for a black-box search.
Appendix F Experimental Setup
F.1 Toy Search Spaces
Below are the hyperparameter settings for the two toy spaces for TangleNAS and SPOS. All experiments were run on a single RTX-2080 GPU.
Search Space Cell Space Conv Macro
Epochs 100 100
Learning rate (LR) 0.1 3e-4
Min. LR 0.001 0.0001
Optimizer SGD Adam
Architecture LR 0.0003 0.0003
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0005 0.0005
Arch Weight Decay 0.001 0.001
Regularization Type L2 L2
Regularization Scale 0.001 0.001
Table 21: Configurations used in the DrNAS experiments on Toy Spaces.
Search Space Cell Space Conv Macro
Epochs 250 250
Learning rate (LR) 0.1 3e-4
Min. LR 0.001 0.0001
Optimizer SGD Adam
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0005 0.0005
Table 22: Configurations used in the SPOS experiments on Toy Spaces.
F.2 Language Model
We use the AdamW optimizer in all experiments related to language modeling. Other hyperparameter choices are as specified in Table 23. Below are the hyperparameter settings for TangleNAS. We run experiments on TinyStories on 8 RTX-2080 GPUs and experiments on OpenWebText on 8 A6000 GPUs.
Search Space Small-LM
Learning rate (LR) 5e-4
Min LR 5e-5
Beta2 0.99
Warmup Iters 100
Max Iters 6000
Lr decay iters 6000
Batch size 12
Weight decay 1e-1
Table 23: Configurations used in DrNAS on the Language Model Spaces.
F.3 AutoFormer and OFA
We use a 50%-50% train and validation split for the CIFAR-10 and CIFAR-100 datasets for the cell-based spaces and a 80%-20% for the weight-entangled spaces. We use the official source code of AutoFormer available at code autoformer for all the AutoFormer experiments on CIFAR-10 and CIFAR-100, closely following the AutoFormer training pipeline and search space design. AutoFormer explored three transformer sizes: Autoformer-T (tiny), AutoFormer-S (small), and Autoformer-B (base). We restrict ourselves to Autoformer-T.
For baselines like OFA and AutoFormer, we follow their respective recipes to obtain the train-validation split for ImageNet-1k. Our models were trained on 2xA100s with the same effective batch-size as AutoFormer. For MobileNetV3 from Once-for-All, we use the same training hyperparameters as the baseline found here (in addition to architectural parameters same as Table 24). We run experiments on CIFAR-10, CIFAR-100, and ImageNet1-k on 4 A100 GPUs.
F.3.1 AutoFormer Fine-tuning
CIFAR-10 and CIFAR-100 pretrained supernet
We fine-tune the CIFAR-10 and CIFAR-100 selected networks (after inheriting them from the supernet) for 500 epochs. We set the learning rate to 1e-3, the warmup epochs to 5, the warmup learning rate to 1e-6, and the minimum learning rate to 1e-5. All other hyperparameters are set the same as in Appendix F.3.
ImageNet pretrained supernet
We follow the DeiT (touvron-icml21) fine-tuning pipeline as used in AutoFormer, to finetune on downstream tasks. Specifically, we set the epochs to 1000, the warmup epochs to 5, the scheduler to cosine, the mixup to 0.8, the smoothing to 0.1, the weight decay to 1e-4, the batch size to 64, the optimizer to SGD, the learning rate to 0.01 and the warmup learning rate to 0.0001 for all datasets. Fine-tuning is performed on 8 RTX-2080 GPUs.
F.4 NB201 and DARTS
For single-stage optimizers, the supernet was trained with four different seeds. The supernet with the best validation performance among these four was discretized to obtain the final model, which was then trained from scratch four times to obtain the results shown in the table. For two-stage methods, we again train the supernet four times and perform Random Search (RS) and Evolutionary Search (ES) on each one. The best model obtained across all four supernets for both methods was then trained from scratch with four seeds to compute the final results. For DrNAS, we follow the same procedure as suggested by the authors across all search spaces. To accommodate multiple training recipes, we have developed a configurable training pipeline. The configurations for DrNAS and SPOS are shown in Tables 24 and 25, respectively. We run our search on a single RTX-2080 for both DARTS and NB201, and train the DARTS architectures from scratch on 8 RTX-2080 GPUs.
Search Space DARTS NB201
Epochs 50 100
Learning rate (LR) 0.1 0.025
Min. LR 0.0 0.001
Architecture LR 0.0006 0.0003
Batch Size 64 64
Momentum 0.9 0.9
Nesterov True False
Weight Decay 0.0003 0.0003
Arch Weight Decay 0.001 0.001
Partial Connection Factor 6 -
Regularization Type L2 L2
Regularization Scale 0.001 0.001
Table 24: Configurations used in the DrNAS experiments.
Search Space DARTS NB201
Epochs 250 250
Learning rate (LR) 0.025 0.025
Min. LR 0.001 0.001
Architecture LR 0.0003 0.0003
Batch Size 256 64
Momentum 0.9 0.9
Nesterov True True
Weight Decay 0.0005 0.0005
Arch Weight Decay 0.001 0.001
Table 25: Configurations used in the SPOS experiments.
Appendix G Optimal architectures derived
G.1 DARTS
Figure 7: DRNAS with WE normal cell
Figure 8: DRNAS with WE reduction cell
G.2 Small LM
num_layers: 7, embed_dim: 768, num_heads: [12, 12, 12, 12, 12, 12, 12], mlp_ratio: [3, 4, 4, 4, 4, 4, 4]
G.3 AutoFormer
G.3.1 CIFAR-10
50%50% split
mlp_ratio:$[4,4,4,4,4,4,4,4,4,3.5,4,4,3.5,4]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, num_layers: 14 embed_dim: 216
80%-20% split
mlp_ratio:$[4,4,4,4,4,4,4,4,3.5,4,4,4,3.5,3.5]$, num_heads:$[4,4,4,4,4,4,4,4,4,4,4,4,4,4]$, depth: 14, embed_dim: 240
G.3.2 CIFAR-100
50%-50% split
mlp_ratio: [4,4,4,4,4,4,4,4,3.5,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4], depth: 14, embed_dim: 216
80%-20% split
mlp_ratio:[3.5,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
G.3.3 ImageNet1-k
mlp_ratio:[4,4,4,4,4,4,4,4,4,4,4,4,4,4], num_heads:[4,4,4,4,4,4,4,4,4,4,4,4,4,4, depth: 14, embed_dim: 240
G.4 MobileNetV3
Kernel_sizes:[7,5,5,7,5,5,7,7,5,7,7,7,5,7,7,5,5,7,7,5],Channel_expansion_factor:[6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6,6], Depths : [4, 4, 4, 4, 4]
G.5 Toy Spaces
G.5.1 Toy cell (our best architecture)
50%-50% split
Genotype(normal=[(’dil_conv_3x3’, 0), (’dil_conv_3x3’, 0), (’sep_conv_3x3’, 1)], normal_concat=range(1, 3), reduce=[(’sep_conv_3x3’, 0), (’sep_conv_3x3’, 0), (’dil_conv_3x3’, 1)], reduce_concat=range(1, 3))
80%-20% split
Genotype(normal=[(’dil_conv_3x3’, 0), (’dil_conv_5x5’, 0), (’sep_conv_3x3’, 1)], normal_concat=range(1, 3), reduce=[(’sep_conv_3x3’, 0), (’sep_conv_5x5’, 0), (’dil_conv_3x3’, 1)], reduce_concat=range(1, 3))
G.6 Toy conv-macro (our best architecture)
50%-50% split:
Channels = [32, 64, 128, 64], Kernel Sizes = [5, 5, 7, 7].
Train-Val fraction 80%-20%:
Channels = [32, 64, 128, 64], Kernel Sizes = [5, 5, 7, 7].
Appendix H Architecture Representation Analysis
CKA
Centered Kernel Alignment (CKA) (kornblith2019similarity) is a metric based on the Hilbert-Schmidt Independence Criterion (HSIC). It is designed to model the similarity between representations in neural networks. In this section, we analyze the CKA between structurally identical layers in the inherited, fine-tuned, and retrained networks in the AutoFormer-T space. Specifically, the aim is to assess how similar the inherited and fine-tuned representations are to those of the models trained from scratch. Table 26 presents the average CKA values for a fixed subset of the CIFAR-10 and CIFAR-100 datasets. We find that the representations learned by the single-stage supernet are more similar to those of the architectures that are fine-tuned or trained from scratch. This observation underscores potential issues with using inherited weights from the supernet as a proxy for search in two-stage methods, as noted by xu2022analyzing.
Model Inherit v/s Retrain Fine-Tune v/s Retrain
CIFAR-10 CIFAR-100 CIFAR-10 CIFAR-100
TangleNAS 0.4630 0.5853 0.5712 0.6527
SPOS+ES 0.4581 0.5793 0.5694 0.6309
SPOS+RS 0.4412 0.583 0.5797 0.6389
Table 26: CKA correlation between layers.
Appendix I Limitations and Future Work
Currently, TangleNAS is designed to optimize a single objective, such as a chosen performance metric. However, in practice, there may be multiple objectives of interest, including hardware efficiency, robustness, and fairness (dooley-neurips23). Additionally, it would be valuable to explore and apply our findings across various applications in computer vision and natural language processing.